OpenStudy (bananas1234):

For the function f(x) = 2 – 3x, find f(–4). –10 –5 12 14

OpenStudy (anonymous):

14

OpenStudy (anonymous):

It would be -10

OpenStudy (anonymous):

14

OpenStudy (bananas1234):

Could you please explain how you got that? @ahamed and @cjcaschera

OpenStudy (anonymous):

Actually on second thought 12 would be a better fit

OpenStudy (bananas1234):

Could you please explain @KatieMTB

OpenStudy (anonymous):

f(-4)= 2-3(-4) = 14 simple as that!

OpenStudy (anonymous):

You just plug the (-4) into the equation 2 - 3x wherever there is an x, so it becomes 2 - 3(-4) and solve.

OpenStudy (anonymous):

I am eating a banana right now!

OpenStudy (bananas1234):

ohhh i see thanks guys!! @ahamed and @KatieMTB

OpenStudy (anonymous):

f(4) = 3*4 = 12

OpenStudy (anonymous):

@cjcaschera it is a (-4)

OpenStudy (bananas1234):

Ha ha thats awesome @ahamed. Gotta love bananas :)

OpenStudy (anonymous):

award me the best answer i was the first one!

OpenStudy (bananas1234):

There @ahamed